# Medium effects on the electromagnetic form factors of $\rho$ meson

### v1

@@ I Introduction
Quantum Chromodynamics (QCD) is widely believed to be an underlying theory of strong interaction of the Standard Model (SM) [1]. The nonperturbative QCD at the low-energy scale depends sensitively on the strong running coupling constant $\alpha_{s}(Q^{2})$. Recent interest in the running coupling has increased in focusing on its behavior in relation to gluon saturation phenomena at small quark longitudinal momentum $x$ (Bjorken-$x$) and the low negative values of the squared four-momentum transfer ($q^{2}\equiv-Q^{2}<0$), where nonlinear QCD effects become significant [2, 3]. Rather different from the perturbative QCD aspects, which are relatively well understood and more straightforward to apply, the nonperturbative QCD, which has the properties of confinement and spontaneous or dynamical chiral symmetry breaking (SCSB or DCSB), is still not fully understood until now. These features of nonperturbative QCD are expected to be better understood by studying the internal structure of hadrons, such as the parton distribution functions (PDFs), electromagnetic form factors (EMFFs), parton distribution amplitudes (PDAs), transverse momentum dependent parton distribution functions (TMD-PDFs), gravitational form factors (GFFs), and generalized parton distribution functions (GPDs), in free space as well as in a nuclear medium. As is well known, hadrons can be described effectively as either three valence dressed quarks for baryons or a dressed valence quark-antiquark pair for mesons, with their internal quark dynamics governed by the strong interaction of QCD.
@@
In the 50 years of the existence of QCD [4], much impressive progress has been made in studying the free space properties of the hadrons to better understand QCD [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. The information on the internal structure of hadrons has been obtained through the PDFs [17, 18], EMFFs [13], PDAs [19, 20, 21], TMDs [22, 23], GFFs [24, 25], and GPDs [26]. These physical properties have been extensively studied, mostly in free space. In addition to the studies in free space, the features of QCD can also be explored by investigating hadrons in a medium, e.g., through their structure functions in nuclei or nuclear matter [27]. The European Muon Collaboration (EMC) experiment provided clear evidence that the hadron unpolarized structure function is modified in a nuclear medium. This modification, known as the EMC effect, is reflected in the ratio of the unpolarized structure function of a bound hadron to that in free space showing less than unity [28], in particular at the large Bjorken-$x$ region. Since then, studies of hadron structure and properties in the nuclear medium have expanded significantly (see Refs. [27, 29, 30] and references therein). This has primarily motivated us to study the $\rho^{+}$ meson EMFFs in the nuclear medium, which constitutes the central focus of the present work.
@@
The $\rho$ mesons have interesting internal quark dynamic features compared to those of the pions. For instance, even though both $\pi^{+}$ and $\rho^{+}$ mesons consist of the $u$ and $\bar{d}$ light dressed quarks and antiquarks, their physical masses, and internal structures are much different because of the spin interactions of the dressed quark and antiquark contents inside the mesons. This difference exhibits that the $\rho$ meson with mass $m_{\rho}=0.770$ GeV has spin1-1 and odd parity $J^{P}=1^{-}$, and the pion with mass $m_{\pi}=0.140$ GeV has spin0-0 and odd parity $J^{P}=0^{-}$. Delving deeper into the $\rho$ meson internal structure will lead us to a better understanding of the contribution of quarks or quark dynamics to its internal structure, and one can gain new insight into physics features based on QCD. So far, to gain a better understanding of the role of the nuclear medium on hadron structure and QCD properties in the medium, several studies have been performed using different theoretical models and approaches in Refs. [21, 31, 32, 33, 34, 35, 36, 37], in particular for the pion EMFFs. This is supported by the availability of experimental data for the free space pion EMFFs, even for only the intermediate $Q^{2}$ range, where $(Q^{2}=-q^{2}>0)$ is the negative of squared four-momentum transfer. However, experimental (empirical) data for the $\rho$ meson EMFFs in free space or in the nuclear medium are still scarce compared to those for the pion. In the past, several experiments have been performed at the CERN SPS through the CERES/NA45 collaboration [38], CLAS g7 at JLAB [39, 40], and KEK Proton-Synchrotron [41, 42] to measure or extract the $\rho$ meson properties in the nuclear medium, such as the mass, the spectral density function, and the decay width. Besides the experimental efforts, extensive theoretical studies have been carried out to investigate the properties of $\rho$ mesons using various models and approaches, such as QCD sum rule (QCDSR) [43, 44], inverse QCD sum rule (IQCDSR) [45], chiral perturbation theory (ChPT) [46], generalized Nambu–Jona-Lasinio (GNJL) model [47], vector meson dominance (VMD) model [48, 49, 50], and dispersion relation [51]. These studies have been performed mostly on the in-medium $\rho$ meson properties focusing on the chiral symmetry restoration, while a study of the $\rho$ meson EMFFs in the nuclear medium is still scarce in the literature.
@@
Recently, investigations of the free-space $\rho$-meson properties, such as its electromagnetic form factors (EMFFs), including the charge radius, magnetic moment, and quadrupole moment, have been extended in order to study their behavior in the nuclear medium [52]. The authors of Ref. [52] reported interesting features of the $\rho$ meson EMFFs in the nuclear medium, which were calculated in a hybrid model; combining the light front constituent quark model (LFCQM) and quark-meson coupling (QMC) model. They found that the quadrupole form factor, $G_{Q}(Q^{2})$, exhibits intriguing behavior in the nuclear medium. In particular, at low momentum transfer ($Q^{2}\simeq 3$ GeV2), the values of $G_{Q}(Q^{2})$ at $\rho_{B}/\rho_{0}=0.25$ and $0.50$ intersect with the free-space result at $\rho_{B}/\rho_{0}=0.0$. This notable feature has motivated us to further investigate the $\rho^{+}$ meson EMFFs in the nuclear medium to better understand the underlying origin and implications of the behavior in the quadrupole form factor.
@@
In the present study, using a consistent treatment approach, we systematically investigate the $\rho^{+}$ meson EMFFs incorporating the charge/electric $G_{C}^{*}(Q^{2})$, the magnetic moment $G_{M}^{*}(Q^{2})$, and the quadrupole moment $G_{Q}^{*}(Q^{2})$ in symmetric nuclear matter within the framework of the covariant NJL model using the Schwinger proper time regularization scheme–simulating the quark confinement, where the medium effect is also calculated within the NJL model at the quark level [53]. The NJL model has been successfully employed in describing the hadron internal structures in free space and nuclear medium, e.g., the $\rho$ meson EMFFs [13, 54], PDF [33], and $\rho$ meson TMD-PDFs [23]. In this work, we compute the $\rho^{+}$ meson effective mass, in-medium $\rho$ meson-quark coupling constant, and in-medium modifications of the $\rho^{+}$ meson EMFFs: charge radius, magnetic moment, and quadrupole moment in the NJL model. As a result, we find that the $\rho$ meson effective mass decreases as the nuclear matter density increases. The reduction in the $\rho$ meson mass at $\rho_{B}/\rho_{0}=1.0$ with $\rho_{0}=0.16$ fm-3 relative to that in free space is about 10%, which is consistent with that obtained by the QCDSR in Ref. [43], and very recent analysis using the IQCDSR in Ref. [45], which are approximated by (10-20)%. However, our result in the $\rho$ mass reduction is rather different from that obtained in the hybrid model of Ref. [52], where they reported the reduction of about 33%. For the $\rho$ meson-quark coupling constant, we find that it decreases as the nuclear matter density increases. In this study, we also find that the in-medium charge radius and quadrupole moment form factors of the $\rho^{+}$ meson increase as the nuclear matter density increases, while the in-medium magnetic moment form factor decreases.
@@ II.1 SU(2) NJL Lagrangian
The NJL model is a Poincaré-covariant quantum field theory-based model that shares QCD’s main properties at low energy scales. For instance, the NJL model encapsulates the emergent features of QCD, such as the dynamical chiral symmetry breaking and confinement in some versions, like the presently used one. The dynamical chiral symmetry is clearly shown in the NJL model through the chiral condensate in the NJL gap equation, as the order parameter of the chiral symmetry breaking, which dynamically generates dressed quark mass from the bare (current) quark mass. As the original NJL model is not a confining model, the confinement in the present study is described through the Schwinger proper-time regularization scheme, which completely removes the unphysical threshold that allows the hadron decay into quarks. An SU(2) flavor symmetric NJL model effective Lagrangian can be written in terms of the four-fermion contact interactions, where the gluons are integrated out and absorbed into the coupling constants with dimension, given by [55]
@@ III NJL model nuclear matter
In this section, we present the extension of the SU(2) flavor NJL model Lagrangian to symmetric nuclear matter, applying the Fierz transformation [56], where the four-fermion interactions can be decomposed as a chiral symmetric linear combination of the form $\sum_{i}G_{i}\big{(}\bar{\psi}_{q}\Gamma_{i}\psi_{q}\big{)}^{2}$. Note that the NJL nuclear matter model is characterized by the nuclear (baryon) density $\rho_{B}$. After performing the quark bilinear  [^1] The quark bilinear $\big{(}\bar{\psi}_{q}\psi_{q}\big{)}$ and $\big{(}\bar{\psi}_{q}\gamma^{\mu}\psi_{q}\big{)}$ can be decomposed as $\big{<}\rho_{B}\mid\bar{\psi}_{q}\psi_{q}\mid\rho_{B}\big{>}+\big{(}:\bar{\psi}_{q}\psi_{q}:\big{)}$ and $\big{<}\rho_{B}\mid\bar{\psi}_{q}\gamma^{\mu}\psi_{q}\mid\rho_{B}\big{>}+\big{(}:\bar{\psi}_{q}\gamma^{\mu}\psi_{q}:\big{)}$. For details, see Ref. [53]. into the effective NJL Lagrangian in Eq. (1), one has [53, 31]
@@
Using the standard hadronization technique [53], the effective potential for symmetric nuclear matter can be determined from the effective NJL Lagrangian. In the mean-field approach (MFA), the energy density of the symmetric nuclear matter is introduced by
@@
$\mathscr{E}_{V}$ $=$ $12i\int\frac{d^{4}k}{(2\pi)^{4}}\ln\Bigg(\frac{k^{2}-M_{q}^{2}+i\epsilon}{k^{2}-M_{0}^{2}+i\epsilon}\Bigg{missing})+\frac{\left(M_{q}-m_{q}\right)^{2}}{4G_{\pi}}-\frac{\left(M_{0}-m_{q}\right)^{2}}{4G_{\pi}},$ (16)
@@ IV $\rho$ meson in nuclear medium
where the asterisk symbol in Eq. (19) stands for the in-medium quantity which enters the light quark momentum $k^{\mu}$ by $k^{*\mu}=k^{\mu}+V^{\mu}$, due to the vector meson mean field $V^{\mu}=\left(V^{0}=V_{0},\mathbf{0}\right)$. Note that the medium modifications of the space component of the light quark momentum $k^{*\mu}$ can be ignored, since the contribution is insignificant [57]. It is to be noted that, after substituting the in-medium modifications of the quark propagator into the constituent quark mass formula in Eq. (18), we can shift the variable of the integral to eliminate the vector potential that enters the light quark momentum.
@@
$t_{\pi}^{*}(p^{*})$ $=$ $\frac{-2iG_{\pi}}{1+2G_{\pi}\Pi_{\pi}^{*}(p^{*2})},~~~~~t_{\rho(\omega)}^{*\mu\nu}(p^{*})=\frac{-2iG_{\rho}}{1+2G_{\rho}\Pi_{V}^{*}(p^{*2}}\Bigg{[}g^{\mu\nu}+2G_{\rho(\omega)}\Pi^{*}_{\rho(\omega)}(p^{*2})\frac{p^{*\mu}p^{*\nu}}{p^{2}}\Bigg{]},$ (20)
@@
$\Pi^{*}_{\pi}(p^{*2})$ $=$ $6i\int\frac{d^{4}k}{(2\pi)^{4}}\mathrm{Tr}\Big{[}\gamma_{5}S_{q}^{*}(k^{*})\gamma_{5}S_{q}^{*}(k^{*}+p^{*}\Big{]},$

### v2

# Medium effects on the electromagnetic form factors of the $\rho$ meson

@@ I Introduction
Quantum chromodynamics (QCD) is widely believed to be an underlying theory of strong interaction of the Standard Model (SM) [Ioffe:2005ym]. The nonperturbative QCD at the low-energy scale depends sensitively on the strong running coupling constant $\alpha_{s}(Q^{2})$. Recent interest in the running coupling has increased in focusing on its behavior in relation to gluon saturation phenomena at small quark longitudinal momentum $x$ (Bjorken-$x$) and the low negative values of the squared four-momentum transfer ($q^{2}\equiv-Q^{2}<0$), where nonlinear QCD effects become significant [Deur:2016tte, Kovchegov:2023bvy]. Rather different from the perturbative QCD aspects, which are relatively well understood and more straightforward to apply, the nonperturbative QCD, which has the properties of confinement and spontaneous or dynamical chiral symmetry breaking (SCSB or DCSB), is still not fully understood until now. These features of nonperturbative QCD are expected to be better understood by studying the internal structure of hadrons, such as the parton distribution functions (PDFs), electromagnetic form factors (EMFFs), parton distribution amplitudes (PDAs), transverse momentum dependent parton distribution functions (TMD-PDFs), gravitational form factors (GFFs), and generalized parton distribution functions (GPDs), in free space as well as in a nuclear medium. As is well known, hadrons can be described effectively as either three valence dressed quarks for baryons or a dressed valence quark-antiquark pair for mesons, with their internal quark dynamics governed by the strong interaction of QCD.
@@
In the 50 years of the existence of QCD [Gross:2022hyw], much impressive progress has been made in studying the free space properties of the hadrons to better understand QCD [Braguta:2004kx, Cardarelli:1994yq, deMelo:2012hj, Jaus:2002sv, Aliev:2009gj, Bhagwat:2006pu, Xu:2024fun, Shi:2023oll, Carrillo-Serrano:2015uca, Qian:2020utg, Haberzettl:2019qpa, Krutov:2018mbu]. The information on the internal structure of hadrons has been obtained through the PDFs [Liu:2023fpj, Kumano:2021xau], EMFFs [Carrillo-Serrano:2015uca], PDAs [Gao:2014bca, Ball:2007zt, Arifi:2023jfe], TMDs [Kumano:2020ijt, Ninomiya:2017ggn], GFFs [Sun:2020wfo, Kim:2022wkc], and GPDs [Sun:2020jng]. These physical properties have been extensively studied, mostly in free space. In addition to the studies in free space, the features of QCD can also be explored by investigating hadrons in a medium, e.g., through their structure functions in nuclei or nuclear matter [Geesaman:1995yd]. The European Muon Collaboration (EMC) experiment provided clear evidence that the hadron unpolarized structure function is modified in a nuclear medium. This modification, known as the EMC effect, is reflected in the ratio of the unpolarized structure function of a bound hadron to that in free space showing less than unity [EuropeanMuon:1983wih], in particular at the large Bjorken-$x$ region. Since then, studies of hadron structure and properties in the nuclear medium have expanded significantly (see Refs. [Geesaman:1995yd, Hayano:2008vn, Hen:2016kwk] and references therein). This has primarily motivated us to study the $\rho^{+}$ meson EMFFs in the nuclear medium, which constitutes the central focus of the present work.
@@
The $\rho$ mesons have interesting internal quark dynamic features compared to those of the pions. For instance, even though both $\pi^{+}$ and $\rho^{+}$ mesons consist of the $u$ and $\bar{d}$ light dressed quarks and antiquarks, their physical masses, and internal structures are much different because of the spin interactions of the dressed quark and antiquark contents inside the mesons. This difference exhibits that the $\rho$ meson with mass $m_{\rho}=0.770$ GeV has spin1-1 and odd parity $J^{P}=1^{-}$, and the pion with mass $m_{\pi}=0.140$ GeV has spin0-0 and odd parity $J^{P}=0^{-}$. Delving deeper into the $\rho$ meson internal structure will lead us to a better understanding of the contribution of quarks or quark dynamics to its internal structure, and one can gain new insight into physics features based on QCD. So far, to gain a better understanding of the role of the nuclear medium on hadron structure and QCD properties in the medium, several studies have been performed using different theoretical models and approaches in Refs. [Arifi:2023jfe, Gifari:2024ssz, Hutauruk:2018qku, Hutauruk:2016sug, Hutauruk:2021kej, Farrar:1979aw, Bernard:1988bx, Hutauruk:2019ipp], in particular for the pion EMFFs. This is supported by the availability of experimental data for the free space pion EMFFs, even for only the intermediate $Q^{2}$ range, where $(Q^{2}=-q^{2}>0)$ is the negative of squared four-momentum transfer. However, experimental (empirical) data for the $\rho$ meson EMFFs in free space or in the nuclear medium are still scarce compared to those for the pion. In the past, several experiments have been performed at the CERN SPS through the CERES/NA45 collaboration [Cassing:1995zv], CLAS g7 at JLAB [CLAS:2008jqp, CLAS:2007dll], and KEK Proton-Synchrotron [Naruki:2005kd, E325:2000smw] to measure or extract the $\rho$ meson properties in the nuclear medium, such as the mass, the spectral density function, and the decay width. Besides the experimental efforts, extensive theoretical studies have been carried out to investigate the properties of $\rho$ mesons using various models and approaches, such as QCD sum rule (QCDSR) [Hatsuda:1991ez, Asakawa:1993pq], inverse QCD sum rule (IQCDSR) [Mutuk:2025lak], chiral perturbation theory (ChPT) [Cabrera:2000dx], generalized Nambu–Jona-Lasinio (GNJL) model [Bernard:1988db], vector meson dominance (VMD) model [Ko:1994hc, Chanfray:1992kp, Rapp:1997fs], and dispersion relation [Kondratyuk:1998ec]. These studies have been performed mostly on the in-medium $\rho$ meson properties focusing on the chiral symmetry restoration, while a study of the $\rho$ meson EMFFs in the nuclear medium is still scarce in the literature.
@@
Recently, investigations of the free-space $\rho$-meson properties, such as its electromagnetic form factors (EMFFs), including the charge radius, magnetic moment, and quadrupole moment, have been extended in order to study their behavior in the nuclear medium [deMelo:2018hfw]. The authors of Ref. [deMelo:2018hfw] reported interesting features of the $\rho$ meson EMFFs in the nuclear medium, which were calculated in a hybrid model; combining the light front constituent quark model (LFCQM) and quark-meson coupling (QMC) model. They found that the quadrupole form factor, $G_{Q}(Q^{2})$, exhibits intriguing behavior in the nuclear medium. In particular, at low momentum transfer ($Q^{2}\simeq 3$ GeV2), the values of $G_{Q}(Q^{2})$ at $\rho_{B}/\rho_{0}=0.25$ and $0.50$ intersect with the free-space result at $\rho_{B}/\rho_{0}=0.0$. This notable feature has motivated us to further investigate the $\rho^{+}$ meson EMFFs in the nuclear medium to better understand the underlying origin and implications of the behavior in the quadrupole form factor.
@@
In the present study, using a consistent treatment approach, we systematically investigate the $\rho^{+}$ meson EMFFs incorporating the charge/electric $G_{C}^{*}(Q^{2})$, the magnetic moment $G_{M}^{*}(Q^{2})$, and the quadrupole moment $G_{Q}^{*}(Q^{2})$ in symmetric nuclear matter within the framework of the covariant NJL model using the Schwinger proper time regularization scheme-simulating the quark confinement, where the medium effect is also calculated within the NJL model at the quark level [Bentz:2001vc]. The NJL model has been successfully employed in describing the hadron internal structures in free space and nuclear medium, e.g., the $\rho$ meson EMFFs [Carrillo-Serrano:2015uca, Cloet:2014rja], PDF [Hutauruk:2016sug], and $\rho$ meson TMD-PDFs [Ninomiya:2017ggn]. In this work, we compute the $\rho^{+}$ meson effective mass, in-medium $\rho$ meson-quark coupling constant, and in-medium modifications of the $\rho^{+}$ meson EMFFs: charge radius, magnetic moment, and quadrupole moment in the NJL model. As a result, we find that the $\rho$ meson effective mass decreases as the nuclear matter density increases. The reduction in the $\rho$ meson mass at $\rho_{B}/\rho_{0}=1.0$ with $\rho_{0}=0.16$ fm-3 relative to that in free space is about 10%, which is consistent with that obtained by the QCDSR in Ref. [Hatsuda:1991ez], and very recent analysis using the IQCDSR in Ref. [Mutuk:2025lak], which are approximated by (10-20)%. However, our result in the $\rho$ mass reduction is rather different from that obtained in the hybrid model of Ref. [deMelo:2018hfw], where they reported the reduction of about 33%. For the $\rho$ meson-quark coupling constant, we find that it decreases as the nuclear matter density increases. In this study, we also find that the in-medium charge radius and quadrupole moment form factors of the $\rho^{+}$ meson increase as the nuclear matter density increases, while the in-medium magnetic moment form factor decreases.
@@ II.1 SU(2) NJL Lagrangian
The NJL model is a Poincaré-covariant quantum field theory-based model that shares QCD’s main properties at low energy scales. For instance, the NJL model encapsulates the emergent features of QCD, such as the dynamical chiral symmetry breaking and confinement in some versions, like the presently used one. The dynamical chiral symmetry is clearly shown in the NJL model through the chiral condensate in the NJL gap equation, as the order parameter of the chiral symmetry breaking, which dynamically generates dressed quark mass from the bare (current) quark mass. As the original NJL model is not a confining model, the confinement in the present study is described through the Schwinger proper-time regularization scheme, which completely removes the unphysical threshold that allows the hadron decay into quarks. An SU(2) flavor symmetric NJL model effective Lagrangian can be written in terms of the four-fermion contact interactions, where the gluons are integrated out and absorbed into the coupling constants with dimension, given by [Klevansky:1992qe]
@@ III NJL model nuclear matter
In this section, we present the extension of the SU(2) flavor NJL model Lagrangian to symmetric nuclear matter, applying the Fierz transformation [Ishii:1993rt], where the four-fermion interactions can be decomposed as a chiral symmetric linear combination of the form $\sum_{i}G_{i}\big(\bar{\psi}_{q}\Gamma_{i}\psi_{q}\big)^{2}$. Note that the NJL nuclear matter model is characterized by the nuclear (baryon) density $\rho_{B}$. After performing the quark bilinear  [^1] The quark bilinear $\big(\bar{\psi}_{q}\psi_{q}\big)$ and $\big(\bar{\psi}_{q}\gamma^{\mu}\psi_{q}\big)$ can be decomposed as $\big<\rho_{B}\mid\bar{\psi}_{q}\psi_{q}\mid\rho_{B}\big>+\big(:\bar{\psi}_{q}\psi_{q}:\big)$ and $\big<\rho_{B}\mid\bar{\psi}_{q}\gamma^{\mu}\psi_{q}\mid\rho_{B}\big>+\big(:\bar{\psi}_{q}\gamma^{\mu}\psi_{q}:\big)$. For details, see Ref. [Bentz:2001vc]. into the effective NJL Lagrangian in Eq. (1), one has [Bentz:2001vc, Gifari:2024ssz]
@@
Using the standard hadronization technique [Bentz:2001vc], the effective potential for symmetric nuclear matter can be determined from the effective NJL Lagrangian. In the mean-field approach (MFA), the energy density of the symmetric nuclear matter is introduced by
@@
$\mathscr{E}_{V}$ $=$ $12i\int\frac{d^{4}k}{(2\pi)^{4}}\ln\Bigg(\frac{k^{2}-M_{q}^{2}+i\epsilon}{k^{2}-M_{0}^{2}+i\epsilon}\Bigg.)+\frac{\left(M_{q}-m_{q}\right)^{2}}{4G_{\pi}}-\frac{\left(M_{0}-m_{q}\right)^{2}}{4G_{\pi}},$ (16)
@@ IV $\rho$ meson in nuclear medium
where the asterisk symbol in Eq. (19) stands for the in-medium quantity which enters the light quark momentum $k^{\mu}$ by $k^{*\mu}=k^{\mu}+V^{\mu}$, due to the vector meson mean field $V^{\mu}=\left(V^{0}=V_{0},\mathbf{0}\right)$. Note that the medium modifications of the space component of the light quark momentum $k^{*\mu}$ can be ignored, since the contribution is insignificant [Krein:1998vc]. It is to be noted that, after substituting the in-medium modifications of the quark propagator into the constituent quark mass formula in Eq. (18), we can shift the variable of the integral to eliminate the vector potential that enters the light quark momentum.
@@
$t_{\pi}^{*}(p^{*})$ $=$ $\frac{-2iG_{\pi}}{1+2G_{\pi}\Pi_{\pi}^{*}(p^{*2})},~~~~~t_{\rho(\omega)}^{*\mu\nu}(p^{*})=\frac{-2iG_{\rho}}{1+2G_{\rho}\Pi_{V}^{*}(p^{*2})}\Bigg[g^{\mu\nu}+2G_{\rho(\omega)}\Pi^{*}_{\rho(\omega)}(p^{*2})\frac{p^{*\mu}p^{*\nu}}{p^{2}}\Bigg],$ (20)
@@
$\Pi^{*}_{\pi}(p^{*2})$ $=$ $6i\int\frac{d^{4}k}{(2\pi)^{4}}\mathrm{Tr}\Big[\gamma_{5}S_{q}^{*}(k^{*})\gamma_{5}S_{q}^{*}(k^{*}+p^{*})\Big],$